# EEmo-Bench: A Benchmark for Multi-modal Large Language Models on Image Evoked Emotion Assessment

### v3

@@ 4.4. Assessment
(3) $\vspace{-2pt}p_{l}=\frac{e^{x_{SCORE\_TOKEN}^{l}}}{\sum_{l}^{l\in\mathcal{L}}e^{x_{SCORE\_TOKEN}^{l}}},$
@@ 5.1. Experiment Setup
In this section, we evaluate the performance of $5$ proprietary MLLMs and $14$ open-source MLLMs, while the latter are divided by parameter scale into two categories: medium-scale (lower than 8b parameters) and large-scale (72b parameters). To be specific, MLLMs are classified into 1) Proprietary MLLMs: Gemini-1.5-pro (Reid et al., 2024), GPT-4o (OpenAI, 2023), Qwen-VL-Plus, Qwen-VL-Max (Bai et al., 2023) and Claude-3.7-Sonnet (Anthropic, 2025). 2) Medium-scale Open-source MLLMs: Deepseek-VL-7B-Chat (Lu et al., 2024), Deepseek-VL2-tiny (Wu et al., 2024a), InternVL2-8B (Team, 2024), InternVL2.5-8B (Chen et al., 2024b), Janus-Pro-7B (Chen et al., 2025), LLaVA-Onevision (Qwen2-7B) (Li et al., 2024a), LLaVA-NeXT (Llama3-8B) (Li et al., 2024b), mPLUG-Owl2 (LLaMA2-7B) (Ye et al., 2024b), mPLUG-Owl3 (Qwen2-7B) (Ye et al., 2024a), Qwen2-VL-7B-Instruct (Wang et al., 2024a), and Qwen2.5-VL-7B-Instruct (Bai et al., 2025). 3) Large-scale Open-source MLLMs: LLaVA-Onevision (Qwen2-72B) (Li et al., 2024a), Qwen2-VL-72B (Wang et al., 2024a), and Qwen2.5-VL-72B (Bai et al., 2025).
@@ 5.2. Observations
Result of Perception Task. The results presented in Tab. 4.3 indicate that all MLLMs significantly surpass the random guessing results in perception tasks, confirming their basic competence in evoked emotions perception. Among these, GPT-4o achieves the highest overall performance at $65.31\%$, closely followed by large-scale open-source models such as LLaVA-Onevision-72B and Qwen2-VL-72B, scoring $64.16\%$ and $63.83\%$, respectively. Notably, large-scale open-source MLLMs generally outperform the proprietary MLLMs (except GPT-4o), underscoring their capacity in image emotion perception tasks. Despite their performance, even the best models fall short of exceptional emotional perception standards, revealing considerable room for improvement in evoked emotions understanding. Detailed observations are provided from the following perspectives.
1) Single vs. Pairs. Results in Tab. 4.3 demonstrate that MLLMs’ average performance for single-image tasks exceeds paired-image tasks by approximately 3.7%, despite a slightly lower random-guess baseline. This suggests that MLLMs have stronger capabilities for emotional perception and reasoning tasks within individual images but face challenges when encountering double-stimulus analysis. (The analysis of the four dimensions, segmented by question concerns, is provided in Supp. A.5.)
2) Perceptual dimensions. Further analysis from Tab. 4.3 reveals that MLLMs effectively recognize salient emotional features, such as valence that derived from colors and representative objects. However, their understanding remains insufficient for more nuanced emotional attributes, like dominance and arousal, that require a deeper integration of human common sense and subjective preferences. This underscores the need to enhance MLLMs’ systematic capabilities regarding emotions evoked by images. (We further illustrate our findings on $6$ content categories in Supp. A.5.)
Result of Ranking Task. As presented in Tab. 4.3, most MLLMs achieve basic performance in ranking the evoked emotions, significantly surpassing random guesses. Among that, Qwen2.5-VL-72B exhibits the highest score at $67.84\%$, closely followed by Qwen-VL-Max and Claude-3.7-Sonnet at $67.27\%$ and $67.05\%$, respectively. However, even these top-performing MLLMs still exhibit a considerable gap compared to human perception. Although capable of identifying primary evoked emotions in many cases, their limited sensitivity to emotional intensity often leads to inconsistent or reversed emotion ranking, ultimately diminishing the overall score. These results underscore the necessity of further enhancing MLLMs’ capabilities in comprehensively perceiving emotional intensity, which is crucial for assessing the variability in emotional experiences across different individuals as well as the diverse emotional responses of a single individual based on images.
Result of Description Task. Based on the results presented in Tab. 5.1, we summarize our key findings as follows:
@@
2) Unbalanced performance across perspectives. MLLMs demonstrate superior performance in comparative tasks involving image pairs but exhibit significant variability across subtasks that focus on different emotional attributes. Specifically, MLLMs demonstrate strong performance on valence, arousal, and reasoning subtasks, with Qwen2.5-VL-72B achieving an excellent performance of $93.87\%$ on valence comparisons, fully meeting daily evaluation requirements for this task. This highlights their strengths in recognizing typical emotional cues derived from symbolic objects and color schemes. Nevertheless, notable shortcomings remain in their nuanced perception and detailed comparison of emotional diversity, particularly regarding dominance. Thus, most MLLMs possess a limited and superficial understanding of image-evoked emotions, underscoring the gap relative to human emotional comprehension.
@@
Result of Assessment Task. From Tab. 3, we observe robust performance for valence, averaging around 0.77, and moderate performance for arousal, averaging approximately 0.46. This finding further confirms their capability in coarse-grained emotional comprehension, particularly regarding valence. However, all models exhibit poor performance in dominance predictions, with the highest correlation (Qwen-VL-7B) reaching only 0.32. Most models yield negatively correlated predictions, likely due to their inability to distinguish clearly between expressed and evoked emotions. For instance, a character with a fierce facial expression and horrible appearance may express a sense of power and dominance but simultaneously evoke feelings of oppression and low dominance in the viewer, causing contradictory outcomes when analyzed by MLLMs predominantly from the perspective of expressed emotions, rather than the supposed standpoint. These observations align closely with the findings from the perception task, further highlighting the limitations of current MLLMs in fine-grained emotional analysis and identifying directions for future improvement in understanding evoked emotional attributes.
@@ 6. Conclusion
In this paper, we introduce EEmo-Bench, the first benchmark for the comprehensive evaluation of the image-evoked emotions perception and understanding capabilities of MLLMs. Our benchmark adopts a carefully curated collection of images from diverse content categories, with rigorous manual annotations, incorporating emotion ranking and the VAD model to assess the diversity and intensity of evoked emotions. Then, a four-dimensional evaluation framework is proposed, including Perception, Ranking, Description, and Assessment. Experimental results from evaluating $19$ MLLMs demonstrate significant variability in model performance across different tasks, with a consistent deficiency in accurately perceiving comprehensive image-evoked emotions and understanding specific emotional attributes. The EEmo-Bench offers critical insights and suggests promising directions for future advancements in the MLLMs IEA field, particularly in enhancing their ability to achieve a thorough and nuanced understanding of image-evoked emotions.
@@ A.1. Details of Annotation Process
The annotation process takes place in a well-controlled laboratory environment to guarantee silence and consistency. Before starting each experiment, we provide a ten-minute break for participants to enter a stable and calm mental state, which is vital for emotional analysis. To ensure the reliability and quality of the annotation, the entire experiment process is divided into $6$ phases, with no more than one and a half hours per phase and no less than half an hour between each phase. The annotation interface is shown in Fig. 4(d) For the assessment of Valence-Arousal-Dominance (VAD), we enumerate some synonyms and a detailed description of each level from the score $1$ to $9$, with five statements of different levels for each attribute. The descriptions are elaborated as follows:
@@
For the description task in EEmo-Bench, we additionally recruited four participants to annotate golden descriptions, with the GUI interface shown in Fig. 4(e). All annotators primarily relied on the emotional attribute labels obtained from the preliminary annotations to answer open-ended questions. This approach minimizes individual subjective emotional biases, thereby simulating the emotional analysis patterns of the majority of people. To ensure the accuracy and rigor of the golden description, each participant is supposed to annotate no more than 30 images a day to avoid fatigue, and every result is meticulously checked and revised by another participant before being adopted.
(d) Annotation GUI for emotion attributes of each image
(e) Annotation GUI for description task of EEmo-Bench
@@ A.2.6. Example Code on Assessment Task
In Algo. $1$, we provide an example code for evaluating the assessment abilities of MLLMs for different attributes. It should be noted that the pseudo code can be easily integrated with any new MLLMs (based on the transformers architecture), enabling them to output quantified predictions of the VAD value.
@@ A.3.5. Evaluation on Description Task
This normalization maps the final score to a range between 00 and $1$. We adopt $S_{ave}$ of each sample for final accuracy representation on the description task. The prompt for evaluation on open-ended questions displays as follows:
@@ A.5. Extended Results for Perception
(1) Question concerns in single images. From Tab. 4, we observe that MLLMs exhibited similar performance in both attributes perception and reasoning perception for single images (except for Gemini-1.5-pro and Claude-3.7-Sonnet), indicating comparable abilities in perceiving emotions and reasoning the causes of emotional responses. Among them, GPT-4o achieved the best performance on both question concerns, with scores of $66.84\%$ and $65.45\%$, respectively, demonstrating its balanced understanding capabilities in single-image emotion analysis.
@@
A.6. Limitation

### v4

@@ 4.4. Assessment
(3) $\vskip-2.0ptp_{l}=\frac{e^{x_{SCORE\_TOKEN}^{l}}}{\sum_{l}^{l\in\mathcal{L}}e^{x_{SCORE\_TOKEN}^{l}}},$
@@ 5.1. Experiment Setup
In this section, we evaluate the performance of 555 proprietary MLLMs and 141414 open-source MLLMs, while the latter are divided by parameter scale into two categories: medium-scale (lower than 8b parameters) and large-scale (72b parameters). To be specific, MLLMs are classified into 1) Proprietary MLLMs: Gemini-1.5-pro (Reid et al., 2024), GPT-4o (OpenAI, 2023), Qwen-VL-Plus, Qwen-VL-Max (Bai et al., 2023) and Claude-3.7-Sonnet (Anthropic, 2025). 2) Medium-scale Open-source MLLMs: Deepseek-VL-7B-Chat (Lu et al., 2024), Deepseek-VL2-tiny (Wu et al., 2024a), InternVL2-8B (Team, 2024), InternVL2.5-8B (Chen et al., 2024a), Janus-Pro-7B (Chen et al., 2025), LLaVA-Onevision (Qwen2-7B) (Li et al., 2024a), LLaVA-NeXT (Llama3-8B) (Li et al., 2024b), mPLUG-Owl2 (LLaMA2-7B) (Ye et al., 2024b), mPLUG-Owl3 (Qwen2-7B) (Ye et al., 2024a), Qwen2-VL-7B-Instruct (Wang et al., 2024a), and Qwen2.5-VL-7B-Instruct (Bai et al., 2025). 3) Large-scale Open-source MLLMs: LLaVA-Onevision (Qwen2-72B) (Li et al., 2024a), Qwen2-VL-72B (Wang et al., 2024a), and Qwen2.5-VL-72B (Bai et al., 2025).
@@ 5.2. Observations
[…]
1) Single vs. Pairs. Results in Tab. 1 demonstrate that MLLMs’ average performance for single-image tasks exceeds paired-image tasks by approximately 3.7%, despite a slightly lower random-guess baseline. This suggests that MLLMs have stronger capabilities for emotional perception and reasoning tasks within individual images but face challenges when encountering double-stimulus analysis. (The analysis of the four dimensions, segmented by question concerns, is provided in Supp. A.5.)
2) Perceptual dimensions. Further analysis from Tab. 1 reveals that MLLMs effectively recognize salient emotional features, such as valence that derived from colors and representative objects. However, their understanding remains insufficient for more nuanced emotional attributes, like dominance and arousal, that require a deeper integration of human common sense and subjective preferences. This underscores the need to enhance MLLMs’ systematic capabilities regarding emotions evoked by images. (We further illustrate our findings on 666 content categories in Supp. A.5.)
[…]
Result of Description Task. Based on the results presented in Tab. 2, we summarize our key findings as follows:
@@
[…]
@@
Result of Assessment Task. From Tab. 3, we observe robust performance for valence, averaging around 0.77, and moderate performance for arousal, averaging approximately 0.46. This finding further confirms their capability in coarse-grained emotional comprehension, particularly regarding valence. However, all models exhibit poor performance in dominance predictions, with the highest correlation (Qwen-VL-7B) reaching only 0.32. Most models yield negatively correlated predictions, likely due to their inability to distinguish clearly between expressed and evoked emotions. For instance, a character with a fierce facial expression and horrible appearance may express a sense of power and dominance but simultaneously evoke feelings of oppression and low dominance in the viewer, causing contradictory outcomes when analyzed by MLLMs predominantly from the perspective of expressed emotions, rather than the supposed standpoint. These observations align closely with the findings from the perception task, further highlighting the limitations of current MLLMs in fine-grained emotional analysis and identifying directions for future improvement in understanding evoked emotional attributes. (We further address failure cases with analysis and improvement suggestions in Supp. A.6.)
@@ 6. Conclusion
In this paper, we introduce EEmo-Bench, the first benchmark for the comprehensive evaluation of the image-evoked emotions perception and understanding capabilities of MLLMs. Our benchmark adopts a carefully curated collection of images from diverse content categories, with rigorous manual annotations, incorporating emotion ranking and the VAD model to assess the diversity and intensity of evoked emotions. Then, a four-dimensional evaluation framework is proposed, including Perception, Ranking, Description, and Assessment. Experimental results from evaluating 191919 MLLMs demonstrate significant variability in model performance across different tasks, with a consistent deficiency in accurately perceiving comprehensive image-evoked emotions and understanding specific emotional attributes. The EEmo-Bench offers critical insights and suggests promising directions for future advancements in the MLLMs IEA field, particularly in enhancing their ability to achieve a thorough and nuanced understanding of image-evoked emotions.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under Grant 62271312 and Grant 62132006, and in part by STCSM under Grant 22DZ2229005.
@@ A.1. Details of Annotation Process
The annotation process takes place in a well-controlled laboratory environment to guarantee silence and consistency. Before starting each experiment, we provide a ten-minute break for participants to enter a stable and calm mental state, which is vital for emotional analysis. To ensure the reliability and quality of the annotation, the entire experiment process is divided into 666 phases, with no more than one and a half hours per phase and no less than half an hour between each phase. The annotation interface is shown in Fig. 5(a) For the assessment of Valence-Arousal-Dominance (VAD), we enumerate some synonyms and a detailed description of each level from the score 111 to 999, with five statements of different levels for each attribute. The descriptions are elaborated as follows:
@@
For the description task in EEmo-Bench, we additionally recruited four participants to annotate golden descriptions, with the GUI interface shown in Fig. 5(b). All annotators primarily relied on the emotional attribute labels obtained from the preliminary annotations to answer open-ended questions. This approach minimizes individual subjective emotional biases, thereby simulating the emotional analysis patterns of the majority of people. To ensure the accuracy and rigor of the golden description, each participant is supposed to annotate no more than 30 images a day to avoid fatigue, and every result is meticulously checked and revised by another participant before being adopted.
To ensure inter-annotator agreement (IAA), we have trained annotators with expertise. The degree of agreement can be quantitatively assessed using the Intraclass Correlation Coefficient (ICC), which is calculated as follows:
(5) $ICC=\frac{MS_{B}-MS_{W}}{MS_{B}+(k-1)MS_{W}+\frac{k}{n}(MS_{R}-MS_{W})},$
where $k$ represents the number of raters, $n$ is the sample size, $MS_{B}$ denotes the between-group variance, $MS_{W}$ is the within-group variance, and $MS_{R}$ refers to the rater variance. ICC values for valence, arousal, and dominance are 0.98850.98850.9885, 0.94140.94140.9414, and 0.97100.97100.9710, respectively, indicating strong consistency between annotators.
(a) Annotation GUI for emotion attributes of each image
(b) Annotation GUI for description task of EEmo-Bench
@@ A.2.6. Example Code on Assessment Task
In Algo. 111, we provide an example code for evaluating the assessment abilities of MLLMs for different attributes. It should be noted that the pseudo code can be easily integrated with any new MLLMs (based on the transformers architecture), enabling them to output quantified predictions of the VAD value.
@@ A.3.5. Evaluation on Description Task
This normalization maps the final score to a range between 0 and 111. We adopt $S_{ave}$ of each sample for final accuracy representation on the description task. The prompt for evaluation on open-ended questions displays as follows:
@@ A.5. Extended Results for Perception
[…]
@@
(a) Arousal analysis.
(b) Dominance analysis.
(c) Perceptual focus misalignment
(d) Insufficient Common Sense Understanding
Figure 10. Typical failure cases of MLLMs on perception and description tasks.
A.6. Observation on Failure Cases
A.6.1. Arousal Analysis
To visually illustrate the difference in MLLMs’ emotional understanding between single images and image pairs, as discussed in Sec. 5.2, we present an analysis of a representative failure case. We evaluate the token limits for all open-source models, ruling out context-window issues and attributing the results to models’ inherent limitations. We find that MLLMs’ understanding of emotional intensity and diversity remains coarse, often recognizing only primary emotions and VAD polarity, preventing nuanced comparisons. As shown in Fig. 10(a), in the comparison of arousal, the left image—featuring a cute puppy with harmonious colors and a shallow depth of field—presents a rich and novel visual experience that stimulates viewer interest. In contrast, the right image, which depicts a surprised mole with an open mouth but relatively muted tones, typically evokes lower arousal level. However, nearly all models provide incorrect predictions, suggesting that MLLMs tend to focus more on the emotional expression of the main subject when interpreting arousal, while remaining less sensitive to visual cues such as color and lighting that can elicit high-arousal emotions like curiosity and surprise.
A.6.2. Dominance Analysis
To support our argument on the performance of dominance assessment discussed in Sec.5.2, we further analyze a failure case illustrated in Fig.10(b). In the left image, a woman shouting conveys a strong expression of dominance but simultaneously induces a sense of oppression in the viewer. In contrast, the right image, featuring a warm smile, expresses moderate dominance yet evokes a heightened sense of psychological dominance, such as warmth and satisfaction. Nevertheless, all models incorrectly classify the left image as eliciting higher dominance.
A.6.3. Improvement Suggestions
Regarding the discussion on enhancing the emotion understanding capabilities of MLLMs in images, we provide a reasoned analysis of the differences in quantitative scores and accuracy in Sec. 5.2. Here, we focus on the discussion of failure cases, providing additional detailed analysis and improvement suggestions as follows.
(1) Three-step perception. After analyzing some failure cases, we find that a three-step COT reasoning approach could help MLLMs better understand the evoked emotions. The process involves: first, identifying typical emotional features (e.g., characters’ expressions); second, perceiving the overall event; and third, incorporating low-level features like color for judgment. For instance, in Fig. 10(c), two children are sitting on an ice rink, waving their arms, with one girl displaying expressions of anger and pain. Despite this, most MLLMs interpret the scene as playful, associating it with positive emotions such as ”joy.” However, a more nuanced interpretation suggests that the children may have collided and blamed each other, which would likely evoke ”anger” in the girl. This example highlights the importance of deeper contextual understanding in the emotional interpretation of scenes.
(2) Commonsense reasoning. MLLMs need to improve their understanding of human common sense and preferences, particularly when it comes to animals. Incorporating social media comments during the pre-training stage of MLLMs could aid in this process. As shown in Supp. A.5 of the paper, even advanced models like GPT-4o and Claude-3.7-Sonnet struggle with interpreting emotions related to animals. For instance, as illustrated in Fig. 10(d), while a close-up of a dragonfly with natural coloring is often linked to positive emotions by MLLMs, it may evoke discomfort for many people, thereby reducing the image’s overall valence. This underscores the importance of understanding people’s inherent impressions and emotional responses to various stimuli, an aspect that MLLMs must learn to interpret more accurately.
A.7. Limitation